# Impurity immersed in a two-component few-fermion mixture in a one-dimensional harmonic trap

### v1

@@ V The structural transition
${\cal F}(g_{0};g_{1},g_{2})=|\langle\mathtt{G}(g_{0},g_{1})|\mathtt{G}(g_{0},g_{2})\rangle|^{2},$ (12)
where we explicitly mark that the many-body ground state $|\mathtt{G}(g_{0},g)\rangle$ is calculated for interactions $g_{0}$ and $g$, respectively. Since for $g_{1}=g_{2}$ the fidelity is extreme and equal to 1, one expands it naturally into the series
${\cal F}_{g_{0}}(g_{1},g_{2})=1-\chi(g_{0};g_{1})(g_{2}-g_{1})^{2}+\ldots$ (13)
In this expansion $\chi(g_{0};g)=\partial^{2}_{g^{\prime}}{\cal F}(g_{0};g,g^{\prime})|_{g^{\prime}=g}$ is the fidelity susceptibility which quantifies resistivity of the ground state on infinitesimal change of interaction strength You et al. (2007); Wang et al. (2015). Its large amplitude for a given value of $g$ signals that at this point the ground state of the system rapidly changes with $g$, i.e., the state of the system quickly moves toward different regions of the Hilbert space.
In Fig. 3(a) we present the susceptibility $\chi(g_{0};g)$ as a function of interaction with the impurity $g$ for different internal interactions $g_{0}$. Successive rows correspond to varying numbers of particles in the mixture $N$. It is clear that the position of the transition $G_{N}(g_{0})$, i.e., interaction strength $g$ for which the susceptibility is maximal (Fig. 3(b)), depends monotonically on the internal interaction in the mixture $g_{0}$. The intensity of the transition (magnitude of the susceptibility) is evidently higher for systems with a larger number of particles. Contrary, the shape of the function $G_{N}(g)$ is not strongly affected by the number of particles.
@@
The observed scaling with the number of particles in the mixture is of course a direct consequence of the scaling of the interaction energy (the impurity interacts mutually with $N^{2}$ particles in the mixture). From this perspective, all susceptibility curves are identical. This also indirectly explains why the scaling found becomes better with an increase in the number of particles.
@@
The numerical fit to all the data (for $N>2$) gives $\Gamma_{0}\approx 4.34$ and the resulting shape is presented in Fig. 3(d) with a dashed black line. The observed universality of the structural transition, and its simple scaling with particle number, suggests that the transition is driven rather by a collective mechanism that is only weakly influenced by the microscopic properties of individual particles.
Lastly, it should be emphasized that the universality just described is a feature for systems where the number of particles in the mixture is large. Indeed, as already mentioned, for small $N$ deviations from the universal curve are evident. Moreover, we checked that for a system of three particles, i.e., a single particle in each component ($N=1$), discrepancies are huge and claiming any universality is not justified.
@@ VI Conclusions
Our work is the next step towards a better understanding of the properties of imbalanced multi-component mixtures of several strongly correlated ultracold fermions. The observed transition and its universality show that a detailed analysis of such systems can expose many of their interesting properties that have no counterparts in systems with fewer components. There are two natural directions for extending our work. One is to consider interactions of the impurity with balanced mixtures of more than two components. The other is to consider interactions with more than one impurity with a multi-component mixture and explore substantial differences with earlier results for a spinless Fermi sea Mistakidis et al. (2019).

### v2

@@ V The structural transition
${\cal F}(g_{0};g,g^{\prime})=|\langle\mathtt{G}(g_{0},g)|\mathtt{G}(g_{0},g^{\prime})\rangle|^{2},$ (13)
where we explicitly mark that the many-body ground state $|\mathtt{G}(g_{0},g)\rangle$ is calculated for interactions $g_{0}$ and $g$, respectively. Since for $g=g^{\prime}$ the fidelity is extreme and equal to 1, one expands it naturally into the series
${\cal F}(g_{0};g,g^{\prime})=1-\chi(g_{0};g)(g^{\prime}-g)^{2}+\ldots$ (14)
In this expansion $\chi(g_{0};g)=\partial^{2}_{g^{\prime}}{\cal F}(g_{0};g,g^{\prime})|_{g^{\prime}=g}$ is the fidelity susceptibility which quantifies the resistivity of the ground state on infinitesimal change of interaction strength You et al. (2007); Wang et al. (2015). Its large amplitude for a given value of $g$ signals that at this point the ground state of the system rapidly changes with $g$, i.e., the state of the system quickly moves toward different regions of the Hilbert space. In our numerical approach, we find susceptibility $\chi(g_{0};g)$ straightforwardly by using a three-point approximation of a second derivative with a small step $\delta g=g^{\prime}-g=10^{-4}$.
In Fig. 5(a), we present the susceptibility $\chi(g_{0};g)$ as a function of interaction with the impurity $g$ for different internal interactions $g_{0}$. Successive rows correspond to varying numbers of particles in the mixture $N$. It is clear that the position of the transition $G_{N}(g_{0})$, i.e., interaction strength $g$ for which the susceptibility is maximal (Fig. 5(b)), depends monotonically on the internal interaction in the mixture $g_{0}$. The intensity of the transition (magnitude of the susceptibility) is evidently higher for systems with a larger number of particles. Contrary, the shape of the function $G_{N}(g)$ is not strongly affected by the number of particles.
@@
The observed scaling with the number of particles in the mixture is the same as the scaling of the internal interaction energy in the two-component mixture. From this perspective, all susceptibility curves are identical. This also indirectly explains why the scaling found becomes better with an increase in the number of particles.
@@
The numerical fit to all the data (for $N>2$) gives $\Gamma_{0}\approx 3.84$ and the resulting shape is presented in Fig. 5(d) with a dashed black line. The observed universality of the structural transition and its simple scaling with particle number suggest that the transition is driven rather by a collective mechanism that is only weakly influenced by the microscopic properties of individual particles.
Lastly, in light of our results, we intend to note that the universality just described is probably a feature for systems where the number of particles in the mixture is large. Indeed, as already mentioned, for smaller $N$, deviations from the universal curve become more pronounced. Moreover, we checked that for the smallest system (three particles, single particle in each component, $N=1$), discrepancies are so huge that claiming any universality is not justified.
@@ VI Conclusions
Our work is the next step towards a better understanding of the properties of imbalanced multi-component mixtures of several strongly correlated ultracold fermions. The observed transition and its universality show that a detailed analysis of such systems can expose many of their interesting properties that have no counterparts in systems with fewer components. There are at least three natural directions for extending our work. One is to consider interactions of the impurity with balanced mixtures of more than two components. The other is to consider interactions with more than one impurity with a multi-component mixture and explore substantial differences with earlier results for a spinless Fermi sea Mistakidis et al. (2019). Thirdly, these directions should be simultaneously extended with an analysis of attractive interactions between all or only certain components of a multi-component system.
@@ Acknowledgments
Appendix A Convergence test
[…]
In our work, to determine the many-body ground state of the Hamiltonian (1), we use the exact diagonalization approach. Its accuracy significantly depends on the energy cut-off $E_{C}$ used during the cropping of the many-body basis. For a given system, we determine a sufficient cut-off by calculating the fidelity between ground states $|\mathtt{G}(E_{C})\rangle$ obtained for successively increased cut-offs (with step $2\hbar\Omega$ due to a mirror symmetry of the system)
$\mathbf{F}(E_{C})=|\langle\mathtt{G}(E_{C})|\mathtt{G}(E_{C}-2\hbar\Omega)\rangle|^{2}.$ (18)
[…]
NN $E_{C}-E_{0}$ $E_{C}$ Threshold ${\cal D}_{\mathrm{Fock}}$
2 24 29 99.8% 35 399
3 20 30 99.0% 53 577
4 22 39 99.0% 186 113
5 24 50 98.0% 520 519
Table 1: Determined energy cut-off $E_{C}$ for different number of particles $N$. All the cut-offs meet the criterion of less than $2\%$ fidelity deviation between successive ground states.